OpenStudy (anonymous):

1 yard=91 cm and Rs.1 = 1.5 cent.The cost of 1 yard of rope is 1 cent,what is the cost of 1 m of rope ?

OpenStudy (anonymous):

Cost of 1 yard= 1 cent Cost of 91 cm =1 cent (because 1 yard=91 cm) So, Cost of 1 cm= 1/91 Cent Cost of 100 cm (i.e 1 m)=100*1/91=100/91 Cent So, cost of 1 m is 100/91 Cent=1.099 Cent To change it in Rs.............1.5 Cent=Rs.1 So, 1 Cent=1/1.5 Cost of 1 m=0.67*1.099=Rs. 0.73